OpenStudy (anonymous):

x (x -7 ) / x ^2 +5x- 24 -5x-27/x^2 +5x -24

OpenStudy (wach):

What are you trying to do with this expression?

OpenStudy (anonymous):

simplification

OpenStudy (lgbasallote):

\[\Large \frac{x(x-7)}{x^2 + 5x - 24} - \frac{5x - 27} {x^2+5x - 24}\] common denominators so just subtract the numerators then copy the denominator \[\Large \implies \frac{x(x-7) - (5x-27)}{x^2 + 5x - 24}\] does that help? or do you need more help?

OpenStudy (anonymous):

ya more help plz

OpenStudy (lgbasallote):

okay...expand the numerator \[\Large \implies \frac{x^2 - 7x - 5x + 27}{x^2 + 5x - 24}\] now combine like terms \[\Large \implies \frac{x^2 - 12x + 27}{x^2 + 5x - 24}\] now factor out the numerator and denominator \[\Large \implies \frac{(x-9)(x-3)}{(x-3)(x+8)}\] do you see the next step?

OpenStudy (anonymous):

(x-9 )/(x+8)

OpenStudy (lgbasallote):

right!

OpenStudy (anonymous):

thanks o lot

OpenStudy (lgbasallote):

welcome